The weights of steers in a herd are distributed normally. The standard deviation is 200lbs and the mean steer weight is 1300lbs. Find the probability that the weight of a randomly selected steer is between 1200 and 1579lbs. Round your answer to four decimal places.
Answer #1

solution:-

given mean = 1300 , standard deviation = 200

p(1200 < x < 1579) = p(1200 - 1300)/200 < z < (1579-1300)/200))

= p(-0.5 < z < 1.4)

= p(z <1.4)? p(z <?0.5)
= 0.9192 - 0.3085
= 0.6107